OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: Using complete sentences, explain how to find the maximum value for each function and determine which function has the largest maximum y-value.

OpenStudy (bibby):

waiting on the graphs

OpenStudy (anonymous):

What's the graph of f(x)?

OpenStudy (anonymous):

f(x) = −4(x − 6)2 + 3

OpenStudy (fortytherapper):

Have you learned about the First Derivative Test?

OpenStudy (solomonzelman):

Normally when you say \(\color{#000000 }{ \displaystyle y=\cos(x) }\) Then, \(\color{#000000 }{ \displaystyle-1<y<1 }\), right? (Because \(\color{#000000 }{ \displaystyle -1<\cos(x)<1 }\))

OpenStudy (solomonzelman):

And in fact, even if you say that \(\color{#000000 }{ \displaystyle y=\cos[a(x)] }\) Then, the cosine is also going to be between -1 and 1

OpenStudy (solomonzelman):

So, the maximum of \(\color{#000000 }{ \displaystyle y=\cos[a(x)] }\) or, \(\color{#000000 }{ \displaystyle y=\cos[x] }\) would be?

OpenStudy (solomonzelman):

Based on your answer to this question, you have to multiply that maximum times 2 (since your cosine function is multiplied times a scale factor of 2), and shift your maximum by 4 units up (since the cosine function you are dealing with is shifted 4 units up).

OpenStudy (solomonzelman):

For \(\color{#000000 }{ \displaystyle -4(x - 6)^2 + 3 }\), it is an opening down parabola, and it's maximum is the vertex (which can be foun by completing the square, or graphing).

OpenStudy (anonymous):

thank you!